An investment had a nominal return of 10.6percent last year. If the real return on investment was only 6.3 percent, what was the inflation rate for the year?

a. 17.57%

b. 3.89%

c. 10.73%

d. 4.49%
Answer #1

(1 + R) = (1 + r)(1 + h)

(1 + 0.106) = (1 + 0.063)(1 + h)

h = [(1 + 0.106)/(1 + 0.063)] - 1

h = 0.0405 or 4.05%